# Beam tracing and profile evolution for localised beams in inhomogeneous plasmas

### v2

@@ III.3.2 Case 2
Suppose that we instead want to generate our Hermite polynomials at the boundary by the following ladder operators:
@@
where $H_{n}(x)H_{m}(y)\exp\left(-\frac{1}{2}\bm{x}\cdot\bm{x}\right)=\hat{\mathcal{L}}_{x}^{n}(\tau=0)\hat{\mathcal{L}}_{y}^{m}(\tau=0)\exp\left(-\frac{1}{2}\bm{x}\cdot\bm{x}\right)$. We have once again defined $\bm{x}(0)=\sqrt{-\mathrm{i}\bm{\Psi}_{w,0}}\bm{w}$.
Converting to dimensionless coordinates $\bm{x}(\tau)=\sqrt{-\mathrm{i}\bm{\Phi}}\bm{w}$, we have:
@@
For $\hat{\mathcal{L}}_{I}$, we simply need to set $\bm{\Phi}(\tau=0)=\bm{\Psi}_{w,0}$ and $\bm{\Psi}(\tau=0)=0$, such that $\bm{\Psi}_{\mathrm{tot}}(\tau=0)=\bm{\Psi}_{w,0}$. This is not always possible, but given that we are only trying to match to boundary conditions, we take $\bm{\Phi}(0)$ to be a diagonalisable symmetric matrix. Given the diagonalised basis of $\bm{\Psi}_{w,0}$, call this $\hat{\bm{e}}_{I}$, we then set:
@@
in order for our ladder operators to reduce to that of the normalised Hermite polynomials.
To see this, note that $\exp\left(-\frac{1}{2}\bm{x}\cdot\bm{x}\right)\exp\left(-\frac{1}{2}\bm{w}\cdot\bm{\Psi}\cdot\bm{w}\right)=\exp\left(-\frac{1}{2}\bm{w}\cdot\bm{\Psi}_{\mathrm{tot}}\cdot\bm{w}\right)$, but $\bm{\Psi}_{\mathrm{tot}}(\tau=0)=\bm{\Psi}_{0}$, so this evolves exactly the same as case 1, and matches the boundary conditions. If one is only interested in the profile created from this process, then one simply evaluates:
$\exp\left(-\frac{\mathrm{i}}{2}\bm{w}\cdot\bm{\Phi}\cdot\bm{w}\right)\hat{\mathcal{L}}_{x}^{n}\hat{\mathcal{L}}_{y}^{m}\exp\left(\frac{\mathrm{i}}{2}\bm{w}\cdot\bm{\Phi}\cdot\bm{w}\right).$ (52)
With this definition for our eigenprofiles of eq. 23, our eigenprofiles have been defined to reduce to the complex Hermite polynomials at the boundary.
@@
Importantly, note that the opposite definitions of $\bm{\Psi}_{\mathrm{tot}}$ and $\bm{\Psi}$ in each case, along with the fact that $\bm{\Phi}(\tau)=\bm{\Psi}_{\mathrm{tot}}-\bm{\Psi}$, means that $\bm{\Phi}(\tau)^{\mathrm{(Case\,1)}}=-\bm{\Phi}(\tau)^{\mathrm{(Case\,2)}}$. This crucial distinction allows our two approaches to function identically, although $\bm{\Phi}$ and $\bm{\Psi}$ are not the same in each case.
Note that the total beam(profile and envelope) must have identical evolution using either approach, as they have the same boundary conditions. This is implicitly built into our construction of the ladder operators, as our evolution equation only admits one possible evolution for any unique boundary condition. The ladder operators commute with the evolution operator $\hat{\mathcal{D}}$, thus if they produce the same boundary profiles, they must evolve identically. Therefore, which approach one chooses ultimately depends on personal preference.
@@
We have to treat the square root of $\bm{\Phi}$ with some nuance. In general, one can take the square-root of any diagonalisable matrix with no issues, save for an ambiguity in sign. It is easy to see that this sign ambiguity will have no effect on our results. However, the square-root of a generic matrix is not defined, thus it may not always be possible to take the square root. However, one should not lose sight of the fact that $\sqrt{\bm{\Phi}}$ is only ever used as a coordinate transformation. The actual expression for our ladder operators do not involve square roots, hence if there is a case in which $\bm{\Phi}$ cannot be square-rooted, direct calculation from the ladder operators is still perfectly feasible and well-defined. In the specific case of Gauss-Hermite beam tracing, we want sensible boundary conditions for incident Gauss-Hermite beams, which need not have a diagonalisable, or square root-able $\bm{\Phi}$. However, we are assured that the real and imaginary parts are separately diagonalisable, and hence we can take the square roots of those. Using that fact, one can always define traditional Hermite polynomial solutions at the boundary if all else fails. For example, one can always set $\bm{\Phi}(0)=\pm 2\mathrm{i}\Im(\bm{\Psi})(0)$, depending on which ladder operator we are using to generate the solutions, and adjust the initial condition for $\bm{\Psi}$ accordingly. One is assured to be able to decompose any boundary electric field into components in the Gauss-Hermite beams with $\bm{x}=\sqrt{\Im(\bm{\Psi})(0)}\bm{w}$ at the boundary. Section V.1.2 goes into more detail for applying this in vacuum.
@@ III.4 Hermite tensors
Assuming that said square root can indeed be taken, we would like to point out the link between our solutions and the Hermite tensorsgrad1949; maheshwari2014. We have essentially derived the coefficient tensor with which we contract our Hermite tensors, where the Hermite tensors are given by:
$\bm{H}^{(N)}=(\bm{x}-\bm{\nabla}_{x})^{N}(1),$ (53)
or similarly with the Gaussian generating function. Before discussing how this relates to our solution, we first want to give a short argument that prescribes the exact expression for Hermite tensors of order N. Firstly, from our ladder operator expression, it is clear that the N-index Hermite tensor must be symmetric in its N indices. Furthermore, if we choose diagonal components of our Hermite tensor, they must reduce to the N-th order Hermite polynomial of the coordinate of that diagonal. That is:
@@
where $_{\{\mu_{1}\mu_{2}...\mu_{N}\}}$’ refers to the symmetrisation operation over the indices. It is then simple to see that applying a rotational transformation to the Hermite tensors actually just rotates the coordinates, as one would expect:
$\bm{H}^{(N)}(\bm{x})\overset{\bm{R}}{\rightarrow}\bm{H}^{(N)}(\bm{R}\bm{x}).$ (57)
In order to obtain explicit polynomials solutions, we have to contract this expression with some coefficient tensor. The coefficient tensor we have then found is just:
@@
Note that in the final expressions, we will not encounter any square-roots, so the expression we obtain is well-defined, as one would expect as our ladder operators did not involve square-roots, and they are only used as mathematical machinery for algebraic manipulation here. This relationship with the Hermite tensors is indeed incredibly useful for our purposes, as it would speed up application of ladder operators for generic boundary conditions, as we have demonstrated that it amounts to contraction of certain vectors, in any order, with the Hermite tensors. For example, if we had to apply some ladder operators N times to match a given boundary condition, we only have to evolve the vectors used to match the boundary condition, then apply them to the N-th order Hermite tensor, which is a constant at any point of evolution.
III.5 Properties of ladder operators
The ladder operators presented here are useful in a general context. They allow us to construct infinite families of solutions to this specific case of the Fokker-Planck equationRisken1996, which bears striking resemblance to the PDE for the multidimensonal Ornstein-Uhlenbeck processVatiwutipong2019. They are thus useful mathematical tools that can be applied in a wider mathematical context for solving similar PDEs in any number of dimensions.
III.5.1 Uniqueness of Solutions
III.5.2 Commutativity
$\mathrm{Comm}]\hat{\mathcal{D}}f=\hat{\mathcal{D}}[\mathrm{Comm}]f.$ (62)
If $\mathrm{Comm}]f=Af$ at the boundary, then it is clear that $\mathrm{Comm}]f=Af$ throughout, given that f is a unique solution. Thus, the commutator of any two ladder operators must remain constant, regardless of what point during the evolution we are at.
We thus only need to evaluate the commutator at a single point, and we will know what it is at all other points. This feature is built into our ladder operators, as when we explicitly evaluate the commutators, we find that:
$\begin{split}[\hat{\mathcal{L}}_{I},\hat{\mathcal{R}}_{J}]&=\bm{u}_{I}(\tau)\cdot\bm{v}_{I}(\tau),\\ &=\bm{u}_{I}(0)\cdot\bm{v}_{J}(0),\\ [\hat{\mathcal{R}}_{I},\hat{\mathcal{L}}_{J}]&=-\bm{u}_{J}(0)\cdot\bm{v}_{I}(0),\\ [\hat{\mathcal{L}}_{I},\hat{\mathcal{L}}_{J}]&=[\hat{\mathcal{R}}_{I},\hat{\mathcal{R}}_{J}]=0.\end{split}$ (63)
Due to this, if $f_{0}$ is the stationary state of $\hat{\mathcal{R}}_{I}$, one can show that:
$\hat{\mathcal{R}}_{J}\hat{\mathcal{L}}_{I}^{n}f_{0}=n(-\bm{u}_{J}(0)\cdot\bm{v}_{I}(0))^{n}\hat{\mathcal{L}}_{I}^{n-1}f_{0},$ (64)
as one would expect from the uniqueness of our solutions. We can trivially perform the same process on the stationary state of $\hat{\mathcal{L}}_{I}$.
III.5.3 Biorthogonality
It turns out that the ladder operators can exhibit biorthogonality. We will subsequently assume that we have chosen a biorthonormal basis at the boundary. We are interested in whether such a basis stays biorthonormal when subject to evolution. We will utilise $\bm{x}=\sqrt{\mathrm{i}\bm{\Phi}^{\mathrm{(Case\,1)}}}\bm{w}$ to define our coordinate system in both cases. It is important to keep track of the fact that this then means that our ladder operators have different forms depending on which case we are applying them to. For case 1, they are:
For case 2 they are:
These definitions we used above are to ensure that the ladder operators produce the same results in each case, asserted via uniqueness. In this coordinate system, the ground states we use to generate our solutions in each case are given by 1 and $\exp\left(-\frac{1}{2}\bm{x}\cdot\bm{x}\right)$ respectively.
This demonstrates that the effect of a ladder operator on case 2 is identical to the effect of the adjoint operator on case 1. Note that when calculated the commutators, we demonstrated earlier that $\bm{u}_{I}\cdot\bm{v}_{J}=\bm{u}_{I}(0)\cdot\bm{v}_{J}(0)$. It then follows that:
$\bm{\mu}_{I}\cdot\bm{\nu}_{J}=\bm{\mu}_{I}(0)\cdot\bm{\nu}_{J}(0),$ (67)
which implies that
We are now in a position to investigate the inner-product of any two states of our system. We will assume that our space is N-dimensional, and that $\bm{u}_{I}(0)\cdot\bm{v}_{J}(0)=\delta_{IJ}$. We keep the case 1 state($g_{n_{1}...n_{i}...n_{N}}$) on the left and the case 2 state($f_{m_{1}...m_{i}...m_{N}}$) on the right, and define the following inner product:
and
Note that although our operators do somewhat function as the inverse or adjoint of each other, they are subject to some scaling, hence we opted to address them is left and right operators.
III.6 Exact Hermite polynomial solutions
@@
With all these properties of our ladder operators established, we will now investigate when the solution will remain as a Hermite polynomial. Since both cases will yield the same result, we will focus on case 1. Assuming that $\bm{x}=\sqrt{\mathrm{i}\bm{\Phi}}\bm{w}$ exists, we see that our ladder operators take the form:
$\begin{split}\hat{\mathcal{L}}_{I}&=\bm{u}_{I}(\tau)\cdot\bm{\nabla}_{w},\\ &=\bm{\mu}_{I}(\tau)\cdot\bm{\nabla}_{x},\\ \hat{\mathcal{R}}_{I}&=\bm{v}_{I}(\tau)\cdot(\bm{w}+\mathrm{i}\bm{\Phi}\bm{\nabla}_{w}),\\ &=\bm{\nu}_{I}(\tau)\cdot(\bm{x}-\bm{\nabla}_{x}).\end{split}$ (74)
We see that the direction for the operators are picked out by $\bm{\mu}_{I}(\tau)$ and $\bm{\nu}_{I}(\tau)$. Since $\bm{\mu}_{I}(\tau)=\sum_{J}\frac{\bm{\mu}_{I}(0)\cdot\bm{\nu}_{J}(0)}{|\bm{\nu}_{J}(\tau)|^{2}}\bm{\nu}_{J}(\tau)$, if we choose a biorthonormal basis for our ladder operators at the plasma boundary, $\bm{\mu}_{I}(\tau)=\frac{\bm{\nu}_{I}(\tau)}{|\bm{\nu}_{I}(\tau)|^{2}}$. Thus, all our ladder operators act in the directions specified by $\bm{\nu}_{I}(\tau)$. However, $\bm{\nu}_{I}(\tau)$ need not be an orthogonal set of vectors, and would not be an orthogonal set of vectors in general. The profile which reduces to the 2D $(n,m)$ Hermite polynomial mode is given by:
@@
If $\bm{v}_{I}(\tau)$ are not orthogonal, then we would not have a single Hermite mode as the solution. Given that we only have a 2D problem, this is still quite tractable. Recall that $\bm{\nu}_{I}=\bm{v}_{I}\sqrt{\mathrm{i}\bm{\Phi}}^{-1}$. Suppose that we have some generic $\bm{\nu}_{x}$ and $\bm{\nu}_{y}$. Choosing $\hat{\bm{e}}_{1}=\hat{\bm{\nu}}_{x}$ to be one of the principal axes for our Hermite polynomials, we then define $\hat{\bm{e}}_{2}=\frac{(1-\hat{\bm{\nu}}_{x}\hat{\bm{\nu}}_{x})\cdot\hat{\bm{\nu}}_{y}}{\sqrt{1-(\hat{\bm{\nu}}_{x}\cdot\hat{\bm{\nu}}_{y})^{2}}}$, which is the normalised projection of $\hat{\bm{\nu}}_{y}$ perpendicular to $\hat{\bm{\nu}}_{x}$. Our ladder operators thus become:
@@
$f_{nm}(\bm{x})=|\bm{\nu}_{x}|^{n}\sum_{k=0}^{m}\left(\left(\begin{array}[]{cc}m\\ k\end{array}\right)(\bm{\nu}_{y}\cdot\hat{\bm{e}}_{1})^{m-k}(\bm{\nu}_{y}\cdot\hat{\bm{e}}_{2})^{k}H_{n+m-k}(\hat{\bm{e}}_{1}\cdot\sqrt{\mathrm{i}\bm{\Phi}}\bm{w})H_{k}(\hat{\bm{e}}_{2}\cdot\sqrt{\mathrm{i}\bm{\Phi}}\bm{w})\right)$ (79)
This quantity is well defined after we expand out everything, as we will not have any square roots of matrices left, in a similar fashion to what happens in eq. 59. We thus see that in general, an $nm$-Hermite mode will evolve into a linear combination of Hermite modes in a rotated basis. Note that this strictly means that unless the directions are orthogonal, it generally cannot be expressed as a simple product of Hermite polynomials. In homogeneous media this is clear cut, as $\bm{\nu}_{x}$ and $\bm{\nu}_{y}$ remain in the same directions throughout the entire evolution, thus, an $nm$-Hermite mode remains a single $nm$-Hermite mode.
At the end of the day, one could always apply the ladder operators directly to evaluate the resulting polynomial, avoiding the need to deal with a sum of products of Hermite polynomials, which somewhat defeats the purpose of having Hermite polynomials. However, note that although a linear combination of Hermite polynomials cannot be expressed as a product of two Hermite polynomials in general, we have not found a way to prove that this is not the case for this particular formula that we obtained, and it may very well be the case that this neatly reduces to a product of two(non-orthogonal) Hermite polynomialsPere1998. In section V.2, we discuss that numerical simulation may be the only way to get some sort of a handle on this issue.

### v3

@@ III.3.2 Case 2
We could instead generate our Hermite polynomials at the boundary by the following ladder operators:
@@
where in a similar fashion to the preceding case, $h_{n}(x)h_{m}(y)=\hat{B}_{x}^{+,n}\hat{B}_{y}^{+,m}\exp\left(-\frac{1}{2}\bm{x}\cdot\bm{x}\right)$. As follows from our earlier definition, the Hermite functions are simply related to the Hermite polynomials by the following:
$h_{n}(x)h_{m}(y)\exp\left(\frac{1}{2}\bm{x}\cdot\bm{x}\right)=H_{n}(x)H_{m}(y)=\exp\left(\frac{1}{2}\bm{x}\cdot\bm{x}\right)\hat{L}_{x}^{+,n}\hat{L}_{y}^{+,m}\exp\left(-\frac{1}{2}\bm{x}\cdot\bm{x}\right).$ (54)
Similarly to the previous case, we convert to dimensionless coordinates $\bm{x}(\tau)=\sqrt{-\mathrm{i}\bm{\Phi}}\cdot\bm{w}$, and we once again define $\bm{x}(0)=\sqrt{-\mathrm{i}\bm{\Psi}_{w,0}}\cdot\bm{w}$. To obtain Hermite functions at the boundary, we once again have to match $\hat{\mathcal{L}}_{I}$ and $\hat{\mathcal{R}}_{I}$ to our previously defined ladder operators in eq. 53:
@@
We simply need to set $\bm{\Phi}(\tau=0)=\bm{\Psi}_{w,0}$ and $\bm{\Psi}(\tau=0)=0$, such that $\bm{\Psi}_{\mathrm{tot}}(\tau=0)=\bm{\Psi}_{w,0}$. Once again, we have to assume that $\bm{\Psi}_{w,0}$ is diagonalisable and that we work in the diagonalised basis of $\bm{\Psi}_{w,0}$. $\Psi_{II,0}$ and $\hat{\bm{e}}_{I}$ are defined identically to case 1, so we then set:
@@
in order for our ladder operators to reduce to that of the normalised elegant Hermite functions. Thus, if $\alpha(\tau)$ is the same Gaussian profile amplitude function that we solved for in eq. 35, the solution given by:
$g_{nm}=\alpha(\tau)\hat{\mathcal{L}}_{x}^{n}\hat{\mathcal{L}}_{y}^{m}\exp\left(\frac{\mathrm{i}}{2}\bm{w}\cdot\bm{\Phi}\cdot\bm{w}\right),$ (57)
reduces to:
$g_{nm}(\tau=0)=H_{n}\left(\sqrt{-\mathrm{i}\Psi_{0,xx}}w_{x}\right)H_{m}\left(\sqrt{-\mathrm{i}\Psi_{0,yy}}w_{y}\right)\alpha(0)\exp\left(\frac{\mathrm{i}}{2}\bm{w}\cdot\bm{\Psi}_{w,0}\cdot\bm{w}\right).$ (58)
Tying everything together, we now show that both choices can be chosen to be consistent with each other and reduce to the same boundary conditions. Note that since $\exp\left(({\mathrm{i}}/{2})\bm{w}\cdot\bm{\Phi}\cdot\bm{w}\right)\exp\left(({\mathrm{i}}/{2})\bm{w}\cdot\bm{\Psi}\cdot\bm{w}\right)=\exp\left(({\mathrm{i}}/{2})\bm{w}\cdot\bm{\Psi}_{\mathrm{tot}}\cdot\bm{w}\right)$, but in both cases, $\bm{\Psi}_{\mathrm{tot}}(\tau=0)=\bm{\Psi}_{0}$, so the solutions given by either approach will be identical. With both cases, the electric field given by the total beam, $\mathcal{P}A^{(0)}\exp(\mathrm{i}\psi)$, match the exact same boundary conditions. If both cases have the same boundary conditions, then by uniqueness of solutions to eq. 24, we are assured that they must be the same throughout the plasma. Thus, we have outlined two cases which have been defined to reduce to the elegant Gauss-Hermite modes at $\tau=0$.
@@
Importantly, note that the opposite definitions of $\bm{\Psi}_{\mathrm{tot}}$ and $\bm{\Psi}$ in each case, along with the fact that $\bm{\Phi}(\tau)=\bm{\Psi}_{\mathrm{tot}}-\bm{\Psi}$, means that $\bm{\Phi}(\tau)^{\mathrm{(Case\,1)}}=-\bm{\Phi}(\tau)^{\mathrm{(Case\,2)}}$. This crucial distinction guarantees that our two approaches yield identical results in both cases. Therefore, which approach one chooses ultimately depends on personal preference.
At this juncture, we will further expound on the ‘ladder’ operation performed by the ladder operators. As we will show in appendix B, the ladder operators as we have just defined actually function as inverses of each other. If we had a solution $f_{nm}=\hat{\mathcal{R}}_{x}^{n}\hat{\mathcal{R}}_{y}^{m}(1)$, then $\hat{\mathcal{L}}_{x}^{i}\hat{\mathcal{L}}_{y}^{j}f_{nm}\propto\hat{\mathcal{R}}_{x}^{n-i}\hat{\mathcal{R}}_{y}^{m-j}(1)=f_{(n-i)(m-j)}$. Thus, we have a ladder of solutions which our operators are capable of lowering and raising us through, which is where the terminology arises from.
@@
In order to analyse the solution at generic $\tau$, not just at $\tau=0$, we have to treat the square root of $\bm{\Phi}$ with some nuance. In general, one can take the square-root of any diagonalisable tensor with no issues, save for an ambiguity in sign. It is easy to see that this sign ambiguity will have no effect on our results. However, the square-root of a generic tensor is not defined, thus it may not always be possible to take the square root. However, one should not lose sight of the fact that $\sqrt{\bm{\Phi}}$ is only ever used as a coordinate transformation. The actual expression for our ladder operators do not involve square roots, hence if there is a case in which $\bm{\Phi}$ cannot be square-rooted, direct calculation from the ladder operators is still perfectly feasible and well-defined. In the specific case of Gauss-Hermite beam tracing, we want sensible boundary conditions for incident Gauss-Hermite beams, for which we only require square root-able $\bm{\Phi}(0)$, which we are ultimately free to specify. For example, we are assured that the real and imaginary parts of a generic $\bm{\Psi}_{w,0}$ are separately diagonalisable, and hence we can take the square roots of those, even if $\bm{\Psi}_{w,0}$ itself is not diagonalisable. Using that fact, one can always define traditional Hermite polynomial solutions at the boundary if all else fails. For example, one can always set $\bm{\Phi}(0)=\pm 2\mathrm{i}\Im(\bm{\Psi}_{w})(0)$, depending on which ladder operator we are using to generate the solutions(‘+’ for case 2 and ‘-’ for case 1), and adjust the initial condition for $\bm{\Psi}_{w}$ accordingly. We are always able to decompose any boundary electric field into components in the Gauss-Hermite beams with $\bm{x}=\sqrt{2\Im(\bm{\Psi}_{w})(0)}\cdot\bm{w}$ at the boundary. Section V.0.2 goes into more detail for applying this in vacuum, where this specific choice has the added advantage of being diagonal in the energy basis. That is, the traditional Hermite modes importantly allow us to decompose the energy into contributions from each Hermite mode, which one is not able to do in the ‘elegant’ Hermite polynomial basis.
@@ III.4 Hermite tensors
Assuming that said square root can indeed be taken, we would like to point out the link between our solutions and the Hermite tensors[grad1949, maheshwari2014]. We have essentially derived the coefficient tensor with which we contract our Hermite tensors, where the Hermite tensors are given by:
$\bm{H}^{(N)}=(\bm{x}-\bm{\nabla}_{x})^{N}(1).$ (59)
A similar expression can be given using the Gaussian generating function, which we do not present here as it has no use to us. Before discussing how this relates to our solution, we first want to give a short argument that prescribes the exact expression for Hermite tensors of order N. Firstly, from our ladder operator expression, it is clear that the N-index Hermite tensor must be symmetric in its N indices. Furthermore, if we choose diagonal components of our Hermite tensor, they must reduce to the N-th order Hermite polynomial of the coordinate of that diagonal. That is:
@@
where $_{\{\mu_{1}\mu_{2}...\mu_{N}\}}$’ refers to the symmetrisation operation over the indices. In order to obtain explicit polynomials solutions, we have to contract this expression with some coefficient tensor. The coefficient tensor we have then found is just:
@@
Note that in the final expressions, we will not encounter any square-roots, so the expression we obtain is well-defined, as one would expect as our ladder operators did not involve square-roots, and they are only used as mathematical machinery for algebraic manipulation here. This relationship with the Hermite tensors is indeed incredibly useful for our purposes, as it would speed up application of ladder operators, as we have demonstrated that it amounts to tensor contraction with the Hermite tensor, thus all one needs to simulate is the evolution of $\bm{v}_{I}$.
III.5 Exact Hermite polynomial solutions
@@
With all these properties of our ladder operators established, we will now investigate the generic evolution of a Hermite polynomial initial condition. Since both cases will yield the same result, we will focus on the approach provided by case 1. Assuming that $\bm{x}=\sqrt{\mathrm{i}\bm{\Phi}}\cdot\bm{w}$ exists, we see that our ladder operators take the form:
$\begin{split}\hat{\mathcal{L}}_{I}&=\bm{u}_{I}(\tau)\cdot\bm{\nabla}_{w},\\ &=\bm{\mu}_{I}(\tau)\cdot\bm{\nabla}_{x},\\ \hat{\mathcal{R}}_{I}&=\bm{v}_{I}(\tau)\cdot(\bm{w}+\mathrm{i}\bm{\Phi}^{-1}\cdot\bm{\nabla}_{w}),\\ &=\bm{\nu}_{I}(\tau)\cdot(\bm{x}-\bm{\nabla}_{x}),\end{split}$ (65)
where $\bm{\mu}_{I}=\bm{u}_{I}(\tau)\cdot\sqrt{\mathrm{i}\bm{\Phi}}$ and $\bm{\nu}_{I}=\bm{v}_{I}(\tau)\cdot\sqrt{\mathrm{i}\bm{\Phi}}^{-1}$ In general, these need not be orthogonal sets of vectors. The profile which reduces to the 2D $(n,m)$ Hermite polynomial mode is given by:
@@
If $\bm{v}_{I}(\tau)$ are not orthogonal, then we would not have a single Hermite mode as the solution. Given that we only have a 2D problem, this is still quite tractable. Recall that $\bm{\nu}_{I}=\bm{v}_{I}\cdot\sqrt{\mathrm{i}\bm{\Phi}}^{-1}$. Suppose that we have some generic $\bm{\nu}_{x}$ and $\bm{\nu}_{y}$. Choosing $\hat{\bm{e}}_{1}=\hat{\bm{\nu}}_{x}$ to be one of the principal axes for our Hermite polynomials, we then define $\hat{\bm{e}}_{2}=\frac{(1-\hat{\bm{\nu}}_{x}\hat{\bm{\nu}}_{x})\cdot\hat{\bm{\nu}}_{y}}{\sqrt{1-(\hat{\bm{\nu}}_{x}\cdot\hat{\bm{\nu}}_{y})^{2}}}$, which is the normalised projection of $\hat{\bm{\nu}}_{y}$ perpendicular to $\hat{\bm{\nu}}_{x}$. Our ladder operators thus become:
@@
$f_{nm}(\bm{x})=|\bm{\nu}_{x}|^{n}\sum_{k=0}^{m}\left(\left(\begin{array}[]{cc}m\\ k\end{array}\right)(\bm{\nu}_{y}\cdot\hat{\bm{e}}_{1})^{m-k}(\bm{\nu}_{y}\cdot\hat{\bm{e}}_{2})^{k}H_{n+m-k}(\hat{\bm{e}}_{1}\cdot\sqrt{\mathrm{i}\bm{\Phi}}\cdot\bm{w})H_{k}(\hat{\bm{e}}_{2}\cdot\sqrt{\mathrm{i}\bm{\Phi}}\cdot\bm{w})\right)$ (70)
This quantity is well defined after we expand out everything, as we will not have any square roots of matrices left, in a similar fashion to what happens in eq. 64. We thus see that in general, an $nm$-Hermite mode will evolve into a linear combination of Hermite modes in a rotated basis. Note that a simple product of Hermite polynomials can always be expressed as a linear combination of Hermite polynomials in some other rotated basis, hence we cannot presently conclude that the solution can only be expressed as a linear combination of Hermite modes. In homogeneous media this is clear cut, as $\bm{\nu}_{x}$ and $\bm{\nu}_{y}$ remain in the same directions throughout the entire evolution, thus, an $nm$-Hermite mode remains a single $nm$-Hermite mode.
At the end of the day, one could always apply the ladder operators directly to evaluate the resulting polynomial, avoiding the need to deal with a sum of products of Hermite polynomials, which somewhat defeats the purpose of having Hermite polynomials. However, note that although a linear combination of Hermite polynomials cannot be expressed as a product of two Hermite polynomials in general, we have not found a way to prove that this is the case for this particular formula that we obtained, and it may very well be the case that this neatly reduces to a product of two(non-orthogonal) Hermite polynomials[Pere1998]. In section VI, we discuss that numerical simulation may be the only way to get some sort of a handle on this issue.